# Magnetized neutral 2SC color superconductivity and possible origin of the inner magnetic field of magnetars

### v1

@@ II Cooper pair condensate at ${\bm{B}}\neq 0$ in the neutral medium
From (37) and (38), one can see that the squared Meissner masses of the charged gluons can be negative in the region $0<\Delta/\delta\mu<\sqrt{2}$ and for the 8th rotated gluon in the region $10<\Delta/\delta\mu<$ 1. The appearance of those negative square Meissner masses is pointing out to the existence of an instability that is called the chromomganetic instability.
III.1 Charged gluons Meissner masses versus baryonic chemical potential at ${\bm{B}}\neq 0$
In Figs. 3 and 4, we plot the ratio of the Meissner mass of charged gluons, $m_{M}^{2}/m_{g}^{2}$, with respect to the so called gluon mass, $m_{g}^{2}=4\alpha_{s}{\bar{\mu}}^{2}/3\pi$, versus the baryonic chemical potentials of interest for NS. Using different parametrizations for the coupling constant amplitude coefficient $\eta$, it is shown in Fig. 3 the results at a moderate magnetic field value of $10^{16}$ G and in Fig. 4 for a higher value of $10^{18}$ G.
In both sets of graphs only when the coupling turns very high (i.e. for $\eta=1.15$) does the Meissner masses stop being tachyionic. Moreover, increasing the coupling strength the graphs move to the left, meaning that the Meissner masses become to be tachyonic at smaller values of the baryonic chemical potential (i.e. at smaller densities).
@@ III.1 Charged gluons Meissner masses versus baryonic chemical potential at ${\bm{B}}\neq 0$
We call attention that the rotated magnetic field also plays a very peculiar additional role on the Meissner masses of the charged quarks in the color superconducting phase. Since the rotated magnetic field contains a component associated with the 8-th gluon, it contributes to the Meissner masses of the charged gluons through the non-Abelian interaction with the charged gluons.
This can be seen taking into account the effective action for the charged gluon fields, $G_{\mu}^{\pm}$, in the background of the in-medium rotated magnetic field, $\tilde{H}$, taken along the z-direction, which is given by Ferrer-Vortices ; MCFL-Vortices
@@
with ${\tilde{\Pi}}_{\mu}=\partial_{\mu}-i{\tilde{e}}{\tilde{A}}_{\mu}$ the covariant derivative, $\tilde{f}_{\mu\nu}=\partial_{\mu}\tilde{A}_{\nu}-\partial_{\nu}\tilde{A}_{\mu}$, the Debye mass, $m_{D}$, the Meissner mass $m_{M}$ and $\lambda$ an arbitrary gauge fixing parameter in the ’t Hooft gauge. Essentially, the effective action (39) is the one of a spin-$1$ charged boson field in a magnetic field. Due to the anomalous magnetic moment term $i{\tilde{e}}{\tilde{f}}_{\mu\nu}G_{\mu}^{+}G_{\nu}^{-}$, even in the case when $m_{M}^{2}>0$, one of the charged gluon modes becomes imaginary once the field surpasses a critical value, i.e., ${\tilde{e}}{\tilde{B}}>{\tilde{e}}{\tilde{B}}_{c}=m_{M}^{2}$. This is the well known ”zero-mode problem” for the Yang-Mills fields in the presence of a magnetic field Ferrer-Vortices -String Theory . The solution to this zero-mode problem is the formation of an inhomogeneous gluon field condensate $\langle G_{\mu}^{\pm}\rangle$ together with the amplification of the applied magnetic field. Following the ansatz employed in Ref. MCFL-Vortices where $\langle G_{1}^{-}\rangle=-i\langle G_{2}^{-}\rangle=G(x,y)$, $\langle G_{3}^{-}\rangle=\langle G_{0}^{-}\rangle=0$ and $\langle G_{i}^{+}\rangle=\langle G_{i}^{-}\rangle^{*}$, the Gibbs free energy reads
@@
where $\tilde{H}$ and $\tilde{B}$ are the in-medium applied and induced magnetic fields respectively, and ${\cal F}_{0}$ is the free energy density of the normal phase ($G=0$) at zero applied field. As can be seen from the third term of the rhs of (40) ${\tilde{e}}\tilde{B}$ enters as a negative square mass. In case that $m_{M}^{2}>0$ if ${\tilde{e}}\tilde{B}>m_{M}^{2}/2$ the corresponding gluon mode becomes tachyonic and the ground state of the system has to be modified. In the case of neutral 2SC matter the situation in most of the region of interest at moderate values of baryonic density is even more pressing since $m_{M}^{2}<0$. Thus, in this case a magnetic field, no matter how weak it can be, will activate the mechanism (i.e. no critical field is needed).
@@
from where we can find the vortex condensate solution $G$ and the induced magnetic field $\tilde{B}$. Note that because of the different sign in the applied and induced field, the field $\tilde{B}$ can be dynamically induced by the contribution of the gluon anomalous magnetic moment $2{\tilde{e}}|G|^{2}$, which comes from the original term in (39) $i{\tilde{e}}{\tilde{f}}_{\mu\nu}G_{\mu}^{+}G_{\nu}^{-}$. Also see that from (42) the induced field $\tilde{B}$ is larger than the applied field $\tilde{H}$. This is the magnetic field boosting mechanism that can be activated in this circumstances. Our goal from now on is to see how this mechanism can work to produce the strong fields exhibit by magnetars.
@@
III.3 Gluon vortices and magnetic field amplification
@@
where $D_{n}(\rho)$ is the parabolic cylinder functions with argument $\rho=\sqrt{2|{\tilde{e}}{\tilde{B}}|}(x^{1}-p^{2}/{\tilde{e}}{\tilde{B}})$ and ${\cal N}=(4\pi|{\tilde{e}}{\tilde{B}}|)^{1/4}/\sqrt{n!}$ is the normalization factor, and $n=0,1,2,\cdots$ denotes the Landau levels. We thus obtain in momentum space
@@
The stability requires that the rest energy of the charged gluons must be smaller than twice the quasiparticles energy gap, i.e.,
@@ IV Possible origin of the inner magnetic field of magnetars
In this section we will discuss a possible mechanism that can serve to generate the inner magnetic field of magnetars. The subclass of neutron stars known as magnetars Thompson , where the anomalous X-ray pulsars (AXP) and the soft gamma-ray repeaters (SGR) are included, exhibits magnetic fields that can reach values of the order of $10^{15}$ G on the star surface Kaspi . Even higher magnetic fields of the order of $10^{17}$ G are expected to be sustained in their interiors, as suggested by various recent observations (see, e.g., Rea and references therein), as well as from theoretical estimations Theor-estimations and according to some numerical simulations Uryu .
One way to explain how it is originated the field amplification that occurs in magnetars the so-called dynamo mechanism Dynamo is considered. This mechanism is responsible for the conversion of kinetic energy of an electrically conducting fluid into magnetic energy. Based on this idea, let us consider the energy equipartition between the star magnetic energy and the rotational kinetic energy which is given by
@@
The magnetic-field strengths found in Fig. 6 are of the order of those observed in millisecond pulsars MillisecondPulsasr , but they are far from the expected $10^{17}$ G inner field expected for magnetars. Hence, we see that the order of magnitude of the magnetic field which corresponds to the macroscopic rotational kinetic energy of the star is not enough to induce the large inner field of the magnetars. This result is suggesting that the origin of this field could be associated with a microscopic mechanism that should be able to boost the inner field to larger values.
We want to propose that precisely the mechanism we are describing in this paper can play this role. From Fig. 5 we see that if the core is formed by neutral 2SC color-superconducting quark matter in the chromagnetic-unstable phase an applied magnetic field of order $10^{15}$ G can be boosted to the expected value of $10^{17}$ G. Remarkably, this boosting mechanism not only works for moderately large applied magnetic field but also for fields with smaller magnitudes let’s say of order $10^{8}$ G.
@@
Thus, if the diquark coupling strength is not so large that makes the squared Meissner mass positive and comparable with $4\Delta^{2}$, the induced magnetic field $\tilde{B}$ becomes of the order of the larger parameter between $\Delta$ and $m_{M}$. Therefore the dependence of the boosted field on the external field exists only implicitly in the gap and Meissner mass, which however are almost constant for small and moderate external fields. Hence, the magnitude of an induced field $\tilde{B}$ for an applied field $\tilde{H}$ of order of $10^{8}\sim 10^{9}$ G will be essentially the same as that for $\tilde{H}$ of order $10^{15}$ G, from where we conclude that in the region with $\mu\in[300,450]$ MeV and $0.751.15>\eta>$ 0.75 the induced field is of order $\tilde{B}\sim 10^{17}$ G as can be seen from Fig. 7 where Eq. (52) was used. There, we see that the expected value of $10^{17}$ G for the inner field of magnetars can be achieved by the proposed mechanism. We call attention that in Fig. 7 the results for $\eta=1.15$ are not plotted. It is because at that coupling $m_{M}^{2}>0$ in the density region under consideration (see Figs. 3 and 4) and we need a strong-field approximation ($\tilde{e}\tilde{B}>m_{M}^{2}/2$) to produce the instability, which is a different situation as the one we are considering in the plots of Fig. 7.
@@ V Concluding remarks
In this paper we attempt to find a mechanism that explains how the magnetar magnetic fields are generated. Magnetars are compact astronomical objects endowed with very large magnetic fields. Using the observational data reported in the McGill Online Magnetar Catalog McGill , we calculated the magnitude of the generated stellar magnetic field by using the correlation between the star rotational energy and its magnetic energy. The found result was not sufficient to reproduce the expected field strength and was suggesting that the star field is not only produced by a dynamo effect of a charged rotational plasma, but for some qualitatively different inner mechanism.
Then, assuming that the star core is formed by neutral 2SC matter at intermediate baryonic densities we showed that at the densities of interest and for diquark coupling moderately large, the phase exhibits the so-called chromomagnetic instability with imaginary Meissner masses for the rotationally charged gluons. In this situation an applied magnetic field of any value will enforce the instability. This is the so called ”zero-mode problem” for spin-one charged fields in the presence of a magnetic field Ferrer-Vortices -String Theory .
To remove the instability in this situation a restructuring of the ground state is proposed by the condensation of gluon vortices. This condensate, in addition of removing the instability, serves to boost the applied magnetic field. We showed that in the parameter region of interest a field of order $10^{8}$ G that could be generated by dynamo effect for the rotational frequencies reported for magnetars McGill , will be boosted to $10^{17}$ G by the mechanism here described. This value of $10^{17}$ G is the one indicated in several works as the most feasible one for the interior of magnetars Theor-estimations . The lowest magnitude observed for the surface magnetic field McGill can be justified taking into account that in going from the core to the surface the magnetic field should decrease in order to satisfy the magnetic flux conservation in such a medium with a very high electric conductivity.
@@ Acknowledgements.
EJF work was supported in part by NSF grant PHY-2013222. B.F. work was supported by the National Natural Science Foundation of China (NSFC) under Grant No. 12075093. EJF is indebted with the Nuclear &\&& Particle Astrophysics group at the Institute of Space Sciences (ICE-CSIC) and specially with Dr. Cristina Manuel, for warm hospitality.

### v2

@@ II Cooper pair condensate at ${\bm{B}}\neq 0$ in the neutral medium
We must also indicate that if the magnetic field is taken into account in the calculation of the polarization operators from where the gluon Meissner masses are obtained, we should expect corrections into Eqs. (37)-(38) proportional to ${\tilde{e}}{\tilde{B}}/{\bar{\mu}}^{2}<<1.$ This is a weak radiative effect that will be of the order of $\alpha_{s}\tilde{e}\tilde{B}$ and that can be neglected as compared with the tree contribution to the Meissner masses of the order of $\tilde{e}\tilde{B}$ produced by the magnetic-field direct interaction with the charged gluons through the anomalous magnetic moment term $i{\tilde{e}}{\tilde{f}}_{\mu\nu}G_{\mu}^{+}G_{\nu}^{-}$, as we will see below. Thus, in our calculations the magnetic field will enter in the Meissner masses (37) and (38) only through the dependence of $\Delta$ and $\mu_{e}$ on this.
From (37) and (38), one can see that the squared Meissner masses of the charged gluons can be negative in the region $0<\Delta/\delta\mu_{e}<\sqrt{2}$ and for the 8th rotated gluon in the region $10<\Delta/\delta\mu_{e}<$ 1. The appearance of those negative square Meissner masses is pointing out to the existence of an instability that is called the chromomagnetic instability.
III.1 Charged gluon Meissner masses versus baryonic chemical potential at ${\bm{B}}\neq 0$
In Figs. 3 and 4, we plot the ratio of the Meissner mass of charged gluons, $m_{M}^{2}/m_{g}^{2}$, with respect to the so-called gluon mass, $m_{g}^{2}=4\alpha_{s}{\bar{\mu}}^{2}/3\pi$, versus the baryonic chemical potentials of interest for NS. Using different parametrizations for the coupling constant amplitude coefficient $\eta$, the results are shown in Fig. 3 at a moderate magnetic-field value of $10^{16}$ G and in Fig. 4 for a higher value of $10^{18}$ G. In both sets of graphs only when the coupling turns very high (i.e. for $\eta=1.15$) does the Meissner masses stop being tachyonic. We call attention to the oscillations in the graphs which are originated by the discontinue change in the occupation of different Landau levels as $\mu$ varies at a constant field. These are the known de Haas-van Alphen oscillations.
@@ III.1 Charged gluon Meissner masses versus baryonic chemical potential at ${\bm{B}}\neq 0$
We should notice that the rotated magnetic field also plays a very peculiar additional role on the Meissner masses of the charged quarks in the color superconducting phase. Since the rotated magnetic field contains a component associated with the 8-th gluon, it contributes to the Meissner masses of the charged gluons through the non-Abelian interaction with the charged gluons.
This can be seen taking into account the effective action for the charged gluon fields, $G_{\mu}^{\pm}$, in the background of the in-medium rotated magnetic field, $\tilde{H}$, taken along the z direction, which is given by Ferrer-Vortices ; MCFL-Vortices
@@
with ${\tilde{\Pi}}_{\mu}=\partial_{\mu}-i{\tilde{e}}{\tilde{A}}_{\mu}$ the covariant derivative, $\tilde{f}_{\mu\nu}=\partial_{\mu}\tilde{A}_{\nu}-\partial_{\nu}\tilde{A}_{\mu}$, the Debye mass, $m_{D}$, the Meissner mass $m_{M}$ and $\lambda$ an arbitrary gauge fixing parameter in the ’t Hooft gauge. Essentially, the effective action (39) is the one of a spin-$1$ charged boson field in a magnetic field. Because of the anomalous magnetic moment term $i{\tilde{e}}{\tilde{f}}_{\mu\nu}G_{\mu}^{+}G_{\nu}^{-}$, even in the case when $m_{M}^{2}>0$, one of the charged gluon modes becomes imaginary once the field surpasses a critical value, i.e., ${\tilde{e}}{\tilde{B}}>{\tilde{e}}{\tilde{B}}_{c}=m_{M}^{2}$. This is the well-known ”zero-mode problem” for the Yang-Mills fields in the presence of a magnetic field Ferrer-Vortices -String Theory . The solution to this zero-mode problem is the formation of an inhomogeneous gluon field condensate $\langle G_{\mu}^{\pm}\rangle$ together with the amplification of the applied magnetic field.
Replacing in (39) with the condensate ansatz employed in Ref. MCFL-Vortices where $\langle G_{1}^{-}\rangle=-i\langle G_{2}^{-}\rangle=G(x,y)$, $\langle G_{3}^{-}\rangle=\langle G_{0}^{-}\rangle=0$ and $\langle G_{i}^{+}\rangle=\langle G_{i}^{-}\rangle^{*}$ we find the system free energy $\cal{F}$. Then, in the presence of an applied rotated magnetic field $\tilde{H}$ and an induced rotated magnetic field $\tilde{B}$ the Gibbs free energy is obtained as ${\cal G}={\cal F}-\tilde{B}\tilde{H}$, which in our case reads
@@
where ${\cal F}_{0}$ is the free energy density of the normal phase ($G=0$) at zero applied field. As can be seen from the third term of the rhs of (40) ${\tilde{e}}\tilde{B}$ enters as a negative square mass. In the case that $m_{M}^{2}>0$ if ${\tilde{e}}\tilde{B}>m_{M}^{2}/2$ the corresponding gluon mode becomes tachyonic and the ground state of the system has to be modified. In the case of neutral 2SC matter the situation in most of the region of interest at moderate values of baryonic density is even more pressing since $m_{M}^{2}<0$. Thus, in this case a magnetic field, no matter how weak it can be, will activate the mechanism (i.e. no critical field is needed).
@@
from where we can find the vortex condensate solution $G$ and the induced magnetic field $\tilde{B}$. Note that because of the different sign in the applied and induced fields, the field $\tilde{B}$ can be dynamically induced by the contribution of the gluon anomalous magnetic moment $2{\tilde{e}}|G|^{2}$, which comes from the original term in (39) $i{\tilde{e}}{\tilde{f}}_{\mu\nu}G_{\mu}^{+}G_{\nu}^{-}$. Also see that from (42) the induced field $\tilde{B}$ is larger than the applied field $\tilde{H}$. This is the magnetic-field boosting mechanism that can be activated in this circumstance. Our goal from now on is to see how this mechanism can work to produce the strong fields exhibited by magnetars.
@@
III.3 Gluon vortices and magnetic-field amplification
@@
where $D_{n}(\rho)$ is the parabolic cylinder functions with argument $\rho=\sqrt{2|{\tilde{e}}{\tilde{B}}|}(x^{1}-p^{2}/{\tilde{e}}{\tilde{B}})$, ${\cal N}=(4\pi|{\tilde{e}}{\tilde{B}}|)^{1/4}/\sqrt{n!}$ is the normalization factor, and $n=0,1,2,\cdots$ denotes the Landau levels. We, thus, obtain in momentum space
@@
The stability requires that the rest energy of the charged gluons must be smaller than twice the quasiparticle energy gap, i.e.,
@@ IV Possible origin of the inner magnetic field of magnetars
In this section we will discuss a possible mechanism that can serve to generate the inner magnetic field of magnetars. The subclass of neutron stars known as magnetars Thompson , where the AXPs and the SGR are included, exhibits magnetic fields that can reach values of the order of $10^{15}$ G on the star surface Kaspi . Even higher magnetic fields of the order of $10^{17}$ G are expected to be sustained in their interiors, as suggested by various recent observations (see, e.g., Rea and references therein), as well as from theoretical estimations Theor-estimations and according to some numerical simulations Uryu .
One way to explain how the field amplification that occurs in magnetars is originated, the so-called dynamo mechanism Dynamo is considered. This mechanism is responsible for the conversion of kinetic energy of an electrically conducting fluid into magnetic energy. Based on this idea, let us consider the energy equipartition between the star magnetic energy and the rotational kinetic energy which is given by
@@
The magnetic-field strengths found in Fig. 6 are of the order of those observed in millisecond pulsars MillisecondPulsasr , but they are far from the expected $10^{17}$ G inner field expected for magnetars. Hence, we see that the order of magnitude of the magnetic field which corresponds to the macroscopic rotational kinetic energy of the star is not enough to induce the large inner field of the magnetars. This result suggests that the origin of this field could be associated with a microscopic mechanism that should be able to boost the inner field to larger values.
We want to propose that precisely the mechanism we are describing in this paper can play this role. From Fig. 5 we see that if the core is formed by neutral 2SC color-superconducting quark matter in the chromomagnetic-unstable phase an applied magnetic field of the order of $10^{15}$ G can be boosted to the expected value of $10^{17}$ G. Remarkably, this boosting mechanism works not only for moderately large applied magnetic field but also for fields with smaller magnitudes, say, of the order of $10^{8}$ G.
@@
Thus, if the diquark coupling strength is not so large that makes the squared Meissner mass positive and comparable with $4\Delta^{2}$, the induced magnetic field $\tilde{B}$ becomes of the order of the larger parameter between $\Delta$ and $m_{M}$. Therefore the dependence of the boosted field on the external field exists only implicitly in the gap and Meissner mass, which, however, are almost constant for small and moderate external fields. Hence, the magnitude of an induced field $\tilde{B}$ for an applied field $\tilde{H}$ of the order of $10^{8}-10^{9}$ G will be essentially the same as that for $\tilde{H}$ of the order of $10^{15}$ G, from where we conclude that in the region with $\mu\in[320,450]$ MeV and $0.751.15>\eta>$ 0.75 the induced field is of the order of $\tilde{B}\sim 10^{17}$ G as can be seen from Fig. 7 where Eq. (52) was used. There, we see that the expected value of $10^{17}$ G for the inner field of magnetars can be achieved by the proposed mechanism. We call attention that in Fig. 7 the results for $\eta=1.15$ are not plotted. It is because at that coupling $m_{M}^{2}>0$ in the density region under consideration (see Figs. 3 and 4) and we need a strong-field approximation ($\tilde{e}\tilde{B}>m_{M}^{2}/2$) to produce the instability, which is a different situation than the one we are considering in the plots in Fig. 7.
@@ V Concluding remarks
In this paper we attempt to find a mechanism that explains how the magnetar magnetic fields are generated. Magnetars are compact astronomical objects endowed with very large magnetic fields. Using the observational data reported in the McGill Online Magnetar Catalog McGill , we calculated the magnitude of the generated stellar magnetic field by using the correlation between the star rotational energy and its magnetic energy. The found result was not sufficient to reproduce the expected field strength and suggests that the star field is not only produced by a dynamo effect of a charged rotational plasma, but for some qualitatively different inner mechanism.
Then, assuming that the star core is formed by neutral 2SC matter at intermediate baryonic densities we showed that at the densities of interest and for diquark coupling moderately large, the phase exhibits the so-called chromomagnetic instability with imaginary Meissner masses for the rotationally charged gluons. In this situation an applied magnetic field of any value will enforce the instability. This is the so-called ”zero-mode problem” for spin-one charged fields in the presence of a magnetic field Ferrer-Vortices -String Theory .
To remove the instability in this situation a restructuring of the ground state is proposed by the condensation of gluon vortices. This condensate, in addition to removing the instability, serves to boost the applied magnetic field. We showed that in the parameter region of interest a field of the order of $10^{8}$ G that could be generated by dynamo effect for the rotational frequencies reported for magnetars McGill , will be boosted to $10^{17}$ G by the mechanism here described. This value of $10^{17}$ G is the one indicated in several works as the most feasible one for the interior of magnetars Theor-estimations . The lowest magnitude observed for the surface magnetic field McGill can be justified taking into account that in going from the core to the surface the magnetic field should decrease in order to satisfy the magnetic flux conservation in such a medium with a very high electric conductivity.
@@ Acknowledgements.
The work of E. J. F. was supported in part by National Science Foundation Grant No. PHY-2013222 and Department of Energy Grant No. DE-SC0022023, the work of A. P, was supported by DOE Grant No. DE-SC0022023 and the work of B. F. was supported by the National Natural Science Foundation of China (NSFC) under Grant No. 12075093. E. J. F. is indebted to the Nuclear and Particle Astrophysics group at the Institute of Space Sciences (ICE-CSIC) and especially to Dr. Cristina Manuel, for warm hospitality.